### v1

@@ 3 Framework
We consider a Bayesian agent who makes a statistical decision. This section describes the agent’s environment, formalizes his conceptual knowledge about that environment, and presents some specific examples.
@@ 3.1 Environment
There is a vector $\theta\equiv(\theta_{1},\ldots,\theta_{K})$ of unknown, real-valued “states.” The agent’s prior on $\theta$ is a probability distribution $\mathbb{P}$ over the $K$-dimensional Euclidean space $\mathbb{R}^{K}$. This distribution is normal with mean $\mu\in\mathbb{R}^{K}$ and variance $\Sigma^{K\times K}$:
@@
The agent observes a sample ${\mathcal{S}}\equiv\{(w^{(i)},y^{(i)})\}_{i=1}^{n}$ of size $n$. Each “observation” $(w^{(i)},y^{(i)})$ comprises a “covariate” $w^{(i)}\in\mathbb{R}^{K}$ with Euclidean length $\lVert w^{(i)}\rVert=1$, and an “outcome”
@@
equal to the sum of $\theta^{T}w^{(i)}$ and an independently normally distributed error $u^{(i)}\sim\mathcal{N}(0,\sigma_{u}^{2})$ with mean zero and variance $\sigma_{u}^{2}>0$.
@@
equal to the mean squared difference between them and the corresponding states.555 Suppose $p_{1},\ldots,p_{K}$ are strictly positive and sum to one. Let $D$ be the $K\times K$ diagonal matrix with ${kk}^{\text{th}}$ entry $Kp_{k}$, and let $a^{\prime}\equiv Da$ and $\theta^{\prime}\equiv D\theta$. Then $L(a^{\prime},\theta^{\prime})=\sum_{k=1}^{K}p_{k}(a_{k}-\theta_{k})^{2}$ is a weighted average of the squared differences between the actions and corresponding states. The weights $p_{1},\ldots,p_{K}$ encode the agent’s preferences: the larger is $p_{k}$, the larger is the loss from taking an action $a_{k}$ different than the state $\theta_{k}$. We focus on the case with $p_{k}=1/K$ for each $k$, which makes $D$ equal the identity matrix and $\theta^{\prime}$ equal $\theta$. However, we can easily generalize our analysis to a setting with non-equal weights by replacing $\theta$ with $\theta^{\prime}$. Then what matters are the eigenvalues and eigenvectors of $\operatorname{\mathbb{V}}(\theta^{\prime})=D\Sigma D^{T}$, rather than those of $\Sigma$. This does not change our results or insights substantively.
Let $\operatorname{\mathbb{E}}$ take expectations with respect to $\mathbb{P}$. The agent chooses the action vector that minimizes his posterior expected loss:666 In Appendix Section A1, we explain how the choice problem (3) is equivalent to a prediction problem that arises in the machine and statistical learning literatures. This equivalence comes from interpreting $\theta_{1},\ldots,\theta_{K}$ as values of an unknown function.
@@ 3.2 Conceptual knowledge
Our definition of conceptual knowledge stems from psychologists’. They define concepts as mental representations that help us describe objects in our environment and how they relate (Murphy,, 2002). Accordingly, our agent’s conceptual knowledge tells him how the states $\theta_{1},\ldots,\theta_{K}$ relate. It allows him to represent
@@
as a linear combination of orthonormal vectors $v_{1},\ldots,v_{K}\in\mathbb{R}^{K}$. These vectors encode structural “features” of $\theta$.777 The vectors $v_{1},\ldots,v_{K}$ may not correspond to physical features of the agent’s environment. Instead they are mental constructs he uses to make sense of his environment. For example, nutrients like nitrogen are mental constructs—no farmer “sees” them. All they can see are the effects of applying fertilizers. This is why we refer to identifying $v_{1},\ldots,v_{K}$ as having “conceptual” knowledge. ,888 In Appendix A1.2, we show how to derive $v_{1},\ldots,v_{K}$ from an approximating “model” that captures the “generalizable structure” of the agent’s environment. The coefficients $\gamma_{1},\ldots,\gamma_{K}\in\mathbb{R}$ encode features’ influences on $\theta$. They are akin to “deep parameters” that determine the “reduced-form” states $\theta_{1},\ldots,\theta_{K}$ via the structural relationships embedded in $v_{1},\ldots,v_{K}$ (Lucas,, 1976).
The agent knows the vectors $v_{1},\ldots,v_{K}$. He does not know the coefficients $\gamma_{1},\ldots,\gamma_{K}$, but he knows some contribute more to the states’ prior variances than others. Specifically, he knows each coefficient $\gamma_{k}$ is independently distributed with variance $\lambda_{k}>0$ non-decreasing in kk.999 It is without loss of generality to assume $\gamma_{1},\ldots,\gamma_{K}$ are independently distributed. This is because $\Lambda$ is positive-semidefinite, and so, by the spectral theorem, there is an orthogonal matrix $A\in\mathbb{R}^{K\times K}$ and diagonal matrix $\Lambda^{\prime}\in\mathbb{R}^{K\times K}$ such that $\Lambda=A\Lambda^{\prime}A^{T}$. Then $V^{\prime}\equiv VA$ is orthogonal and $\Sigma$ has eigendecomposition $V^{\prime}\Lambda^{\prime}(V^{\prime})^{T}$, so we can carry out our analysis by replacing $V$ with $V^{\prime}$ and $\Lambda$ with $\Lambda^{\prime}$. Likewise, it is without loss to assume $\lambda_{1}\geq\cdots\geq\lambda_{K}$ because we can permute the indices of the eigenpairs $(\lambda_{k},v_{k})$ without changing $\Sigma$. Then $\theta$ has prior variance
@@
is the $K\times K$ diagonal matrix with entries $\lambda_{1},\ldots,\lambda_{K}$ and
@@
The distribution of $\lambda_{1},\ldots,\lambda_{K}$ around their mean $\overline{\lambda}=\operatorname{tr}(\Sigma)/K$ captures the states’ “reducibility”: they are more “reducible” when they are influenced by fewer features, which happens precisely when $\lambda_{1},\ldots,\lambda_{K}$ are more spread out around $\overline{\lambda}$.101010 We formalize what it means for $\lambda_{1},\ldots,\lambda_{K}$ to be “more spread out” in Section 4.5. ,111111 If $\lambda_{2}=\cdots=\lambda_{K}$ (as in Example 1), then the distribution of $\lambda_{1},\ldots,\lambda_{K}$ is fully determined by the leading eigenvalue $\lambda_{1}$ and the “spectral gap” $(\lambda_{1}-\lambda_{2})$. This gap appears elsewhere in the statistical literature: it determines Markov chains’ mixing times (Levin et al.,, 2008) and whether principal components can be estimated consistently (Yu et al.,, 2015). Thus, the agent’s conceptual knowledge allows him to “reduce” states by representing them as low-dimensional combinations of higher-dimensional features.
If the agent had no conceptual knowledge, then he would not be able to reduce states because he would not know how to represent them as combinations of features with different influences. So his prior variance matrix
@@
would equal the prior variance matrix in the case when $\lambda_{k}=\overline{\lambda}$ for each $k\in\{1,\ldots,K\}$.121212 The naïve prior variance matrix $\Sigma^{(0)}$ is robust to mis-specification in that it commits as little as possible to any given covariance structure. By spreading variance evenly across all dimensions of $\mathbb{R}^{K}$, it avoids overweighting components of $\theta$ that later prove irrelevant. However, this robustness comes at a cost: it forfeits any gains that could be obtained by overweighting components of $\theta$ that later prove essential.
We use the matrix $\Sigma^{(0)}$ to measure how much the agent’s conceptual knowledge allows him to reduce states. We consider a counterfactual “naïve” agent whose prior $\mathbb{P}^{(0)}\equiv\mathcal{N}(\mu,\Sigma^{(0)})$ on $\theta$ has a different variance than the “true” agent’s prior $\mathbb{P}=\mathcal{N}(\mu,\Sigma)$. The naïve agent does not how the states relate structurally and, thus, assumes $\theta$ has prior variance $\Sigma^{(0)}$. The true agent knows how the states relate structurally and, thus, knows $\theta$ has prior variance $\Sigma$.
Since $\mathbb{P}$ and $\mathbb{P}^{(0)}$ are normal distributions with equal means, the Kullback-Leibler (hereafter “KL”) divergence from $\mathbb{P}$ and $\mathbb{P}^{(0)}$ equals131313 See Rasmussen and Williams, (2006, Section A.5) for a derivation of (6).
@@
The KL divergence (6) measures the “information gain” from using $\mathbb{P}$ as a prior rather than $\mathbb{P}^{(0)}$. This information is purely conceptual: it does not depend on the sample $\mathcal{S}$}. It comes from knowing how to represent states as low-dimensional combinations of high-dimensional features. We study the value of this dimension reduction in Section 5.
@@
The KL divergence (6) depends on the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ of the “true” prior variance matrix $\Sigma$. It equals zero when $\lambda_{1},\ldots,\lambda_{K}$ are equal (to $\overline{\lambda}$) and is larger when they are more spread out (see Proposition A4).141414 If $\mathbb{P}^{(0)}$ has mean $\mu^{(0)}\in\mathbb{R}^{K}$, then (6) becomes $\mathcal{D}_{\text{KL}}(\mathbb{P}\parallel\mathbb{P}^{(0)})=\frac{1}{2}\mathopen{}\mathclose{{\left(\frac{1}{\overline{\lambda}}\lVert\mu-\mu^{(0)}\rVert-\sum_{k=1}^{K}\ln\mathopen{}\mathclose{{\left(\frac{\lambda_{k}}{\overline{\lambda}}}}\right)}}\right).$ So even if $\mu^{(0)}\not=\mu$, the KL divergence from $\mathbb{P}$ to $\mathbb{P}^{(0)}$ is non-negative and does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS (see Proposition A4). But it is strictly larger than zero when $\mu^{(0)}\not=\mu$, even if $\lambda_{1}=\cdots=\lambda_{K}$. If some features are more influential than others, then identifying the most influential features empowers the agent to reduce $\theta_{1},\ldots,\theta_{K}$. Intuitively, identifying these features tells him “where to look” for the vector $\theta$ in the space $\mathbb{R}^{K}$ containing it.

### v2

@@ 3 Framework
We consider a Bayesian agent who collects data before making a statistical decision. This section describes the agent’s environment, formalizes his conceptual knowledge about that environment, explains our modeling assumptions, and presents some specific examples.
@@ 3.1 Environment
There is a true but unknown vector $\theta\equiv(\theta_{1},\ldots,\theta_{K})$ of real-valued “states.” The agent’s prior on $\theta$ is a probability distribution $\mathbb{P}$ over the $K$-dimensional Euclidean space $\mathbb{R}^{K}$. This distribution is normal with mean $\mu\in\mathbb{R}^{K}$ and variance $\Sigma\in\mathbb{R}^{K\times K}$:
@@
The agent derives $\mathbb{P}$ from his conceptual knowledge about $\theta$. We explain this derivation in Section 3.2.
@@
The agent observes a sample ${\mathcal{S}}\equiv\{(w^{(i)},y^{(i)})\}_{i=1}^{n}$ of size $n$. Each “observation” $(w^{(i)},y^{(i)})$ comprises a “covariate” $w^{(i)}\in\mathbb{R}^{K}$ with Euclidean length $\lVert w^{(i)}\rVert=1$, [^8] Assuming the covariates have unit length normalizes the scales of the signals $y^{(1)},\ldots,y^{(n)}$ so that only the directions of $w^{(1)},\ldots,w^{(n)}$ (and not their magnitudes) affect signals’ informativeness. It also ensures the Gram matrix (14) always has trace $n$ (see Footnote 25). This allows us to identify optimal samples of size $n$ via a specific Gram matrix (22). and an “outcome”
@@
equal to the sum of
$\theta^{T}w^{(i)}=\sum_{k=1}^{K}\theta_{k}w_{k}^{(i)}$
and an independently normally distributed error $u^{(i)}$ with mean zero and variance $\sigma_{u}^{2}>0$. Thus, the outcome $y^{(i)}$ provides a noisy signal of a weighted combination of states, where the weights are determined by the covariate $w^{(i)}\equiv(w_{1}^{(i)},\ldots,w_{K}^{(i)})$.
@@
equal to the mean squared difference between them and the corresponding states. [^9] Suppose $p_{1},\ldots,p_{K}$ are strictly positive and sum to one. Let $D$ be the $K\times K$ diagonal matrix with ${kk}^{\text{th}}$ entry $Kp_{k}$, and let $a^{\prime}\equiv Da$ and $\theta^{\prime}\equiv D\theta$. Then $L(a^{\prime},\theta^{\prime})=\sum_{k=1}^{K}p_{k}(a_{k}-\theta_{k})^{2}$ is a weighted average of the squared differences between the actions and corresponding states. The weights $p_{1},\ldots,p_{K}$ encode the agent’s preferences: the larger is $p_{k}$, the larger is the loss from taking an action $a_{k}$ different than the state $\theta_{k}$. We focus on the case with $p_{k}=1/K$ for each $k$, which makes $D$ equal the identity matrix and $\theta^{\prime}$ equal $\theta$. However, we can easily generalize our analysis to a setting with non-equal weights by replacing $\theta$ with $\theta^{\prime}$. Then what matters are the eigenvalues and eigenvectors of $\operatorname{\mathbb{V}}(\theta^{\prime})=D\Sigma D^{T}$, rather than those of $\Sigma$. This does not change our results or insights substantively.
Let $\operatorname{\mathbb{E}}$ take expectations with respect to the prior distribution $\mathbb{P}$. The agent chooses the action vector that minimizes his posterior expected loss: [^10] In Appendix Section A1, we explain how the choice problem (3) is equivalent to a prediction problem that arises in the machine and statistical learning literatures. This equivalence comes from interpreting $\theta_{1},\ldots,\theta_{K}$ as values of an unknown function.
@@
Intuitively, he wants to estimate $\theta_{1},\ldots,\theta_{K}$ accurately, and the accuracy of his estimates $a_{1},\ldots,a_{K}$ is determined by their squared errors $(\theta_{k}-a_{k})^{2}$. Thus, our framework aligns with least squares estimation, a tool used throughout empirical economics and statistics. [^11] Indeed, if the prior distribution $\mathbb{P}$ is diffuse, then (3) equals the Ordinary Least Squares estimate of $\theta$ given $\mathcal{S}$}.
Value of $\mathcal{S}$}
If the agent did not observe the sample $\mathcal{S}$}, then his minimized prior and posterior expected losses would be equal. The information in $\mathcal{S}$} is instrumentally valuable because it helps the agent take actions with lower expected losses. Accordingly, we define the “value of $\mathcal{S}$}” to be the difference between his minimized prior and posterior expected losses: [^12] Raiffa and Schlaifer, (1961, p. 90) define a similar object and call it the “(expected) value of sample information.”
$\pi({\mathcal{S}})\equiv\min_{a^{\prime}\in\mathbb{R}^{K}}\operatorname{\mathbb{E}}[L(\theta,a^{\prime})]-\min_{a^{\prime}\in\mathbb{R}^{K}}\operatorname{\mathbb{E}}[L(\theta,a^{\prime})\mid{\mathcal{S}}].$ (4)
Covariate selection
The agent chooses the covariates $w^{(1)},\ldots,w^{(n)}$ that maximize (4) subject to the length constraints $\lVert w^{(i)}\rVert=1$. Intuitively, he wants to design the signals $y^{(1)},\ldots,y^{(n)}$ so that they provide as much payoff-relevant information as possible about the state vector $\theta$.
Timing
First, nature draws $\theta$ from the prior distribution $\mathbb{P}$. Second, the agent chooses the covariates $w^{(1)},\ldots,w^{(n)}$ and observes the outcomes $y^{(1)},\ldots,y^{(n)}$. Third, he combines $\mathbb{P}$ and the sample ${\mathcal{S}}\equiv\{(w^{(i)},y^{(i)})\}_{i=1}^{n}$ to form posterior beliefs about $\theta$. Finally, he chooses the action vector (3) that minimizes his posterior expected loss.
@@ 3.2 Conceptual knowledge
Mental model and concepts
Our definition of conceptual knowledge draws upon psychologists’ and cognitive scientists’: concepts are the building blocks of mental models (Johnson-Laird,, 1983), are used to describe objects and how they relate (Murphy,, 2002), and allow humans to generalize across objects (Mitchell,, 2021).
Accordingly, our agent’s conceptual knowledge allows him to describe the states $\theta_{1},\ldots,\theta_{K}$ and how they relate. It gives him a mental model of
@@
as an unknown combination of known vectors $v_{1},\ldots,v_{K}\in\mathbb{R}^{K}$. We call these vectors “concepts.” [^13] The vectors $v_{1},\ldots,v_{K}$ may not correspond to physical features of the agent’s environment. Instead they are mental constructs he uses to make sense of his environment. For example, nutrients like nitrogen are mental constructs: no farmer sees them. All farmers see are the effects of applying fertilizers. This is why we call $v_{1},\ldots,v_{K}$ “concepts.” They are the building blocks of the agent’s mental model. They capture his environment’s generalizable structure: each state $\theta_{j}$ depends on the $j^{\text{th}}$ component of $v_{k}$ via a coefficient $\gamma_{k}\in\mathbb{R}$ that is independent of jj. [^14] The coefficients $\gamma_{1},\ldots,\gamma_{K}$ are akin to “deep parameters” that determine the “reduced-form” states $\theta_{1},\ldots,\theta_{K}$ via the structural relationships encoded by $v_{1},\ldots,v_{K}$ (Lucas,, 1976). This allows the agent to generalize across states: signals of $\theta_{1}$ provide information about $\gamma_{1},\ldots,\gamma_{K}$, from which he can extrapolate $\theta_{2},\ldots,\theta_{K}$.
For example, suppose $\theta_{1},\ldots,\theta_{K}$ are the effects of applying different fertilizers. Then $v_{1},\ldots,v_{K}$ could encode nutrient quantities and $\gamma_{1},\ldots,\gamma_{K}$ the fertilizer-invariant effects of supplying different nutrients. [^15] Alternatively, if $\theta_{1},\ldots,\theta_{K}$ are the prices of financial assets, then $v_{1},\ldots,v_{K}$ could encode payoffs in different states of nature and $\gamma_{1},\ldots,\gamma_{K}$ the prices of Arrow-Debreu securities (see, e.g., Varian,, 1987). Moreover, if a farmer learns one fertilizer’s overall effect, then he can extrapolate the others’ via their joint dependence on $\gamma_{1},\ldots,\gamma_{K}$.
For convenience and without loss of generality, we assume $v_{1},\ldots,v_{K}$ are orthonormal for the remainder of the paper.
Eigendecomposition
The agent knows the concepts $v_{1},\ldots,v_{K}$ included in his mental model (5). He does not know the coefficients $\gamma_{1},\ldots,\gamma_{K}$, but he knows some contribute more to the states’ prior variances than others. Specifically, he knows each coefficient $\gamma_{k}$ is independently distributed with variance $\lambda_{k}>0$ non-decreasing in kk. [^16] It is without loss of generality to assume $\gamma_{1},\ldots,\gamma_{K}$ are independently distributed. This is because $\Lambda$ is positive-semidefinite, and so, by the spectral theorem, there is an orthogonal matrix $A\in\mathbb{R}^{K\times K}$ and diagonal matrix $\Lambda^{\prime}\in\mathbb{R}^{K\times K}$ such that $\Lambda=A\Lambda^{\prime}A^{T}$. Then $V^{\prime}\equiv VA$ is orthogonal and $\Sigma$ has eigendecomposition $V^{\prime}\Lambda^{\prime}(V^{\prime})^{T}$, so we can carry out our analysis by replacing $V$ with $V^{\prime}$ and $\Lambda$ with $\Lambda^{\prime}$. Likewise, it is without loss to assume $\lambda_{1}\geq\cdots\geq\lambda_{K}$ because we can permute the indices of the eigenpairs $(\lambda_{k},v_{k})$ without changing $\Sigma$. Then $\theta$ has prior variance
@@
is the $K\times K$ diagonal matrix with entries $\lambda_{1}\geq\cdots\geq\lambda_{K}\geq 0$ and
@@
Reducibility
The distribution of $\lambda_{1},\ldots,\lambda_{K}$ around their mean $\overline{\lambda}=\operatorname{tr}(\Sigma)/K$ captures the states’ “reducibility.” They are more “reducible” when their prior variances are explained by fewer common concepts: when $\lambda_{1},\ldots,\lambda_{K}$ are more spread out around $\overline{\lambda}$. [^17] We formalize what it means for $\lambda_{1},\ldots,\lambda_{K}$ to be “more spread out” in Section 4.4. , [^18] If $\lambda_{2}=\cdots=\lambda_{K}$ (as in Example 1), then the distribution of $\lambda_{1},\ldots,\lambda_{K}$ is fully determined by the leading eigenvalue $\lambda_{1}$ and the “spectral gap” $(\lambda_{1}-\lambda_{2})$. This gap appears elsewhere in the statistics literature; for example, spectral gaps determine Markov chains’ mixing times (Levin et al.,, 2008) and whether principal components can be estimated consistently (Yu et al.,, 2015). The agent’s conceptual knowledge allows him to “reduce” the state vector $\theta$ by representing it as a low-dimensional combination of higher-dimensional concepts.
If the agent had no conceptual knowledge—i.e., if he did not have a mental model of $\theta$ as a combination of concepts with different explanatory powers—then he would not be able to reduce states in the manner described above. His prior variance matrix
@@
would equal the prior variance matrix in the case when $\lambda_{k}=\overline{\lambda}$ for each $k\in\{1,\ldots,K\}$. We refer to $\mathbb{P}^{(0)}\equiv\mathcal{N}(\mu,\Sigma^{(0)})$ as a “naïve” prior because it ignores the covariances among states stemming from their dependence on common concepts.
We can use the true prior $\mathbb{P}\equiv\mathcal{N}(\mu,\Sigma)$ and naïve prior $\mathbb{P}^{(0)}$ to measure how much the agent’s conceptual knowledge allows him to reduce states. Since $\mathbb{P}$ and $\mathbb{P}^{(0)}$ are normal distributions with equal means, the Kullback-Leibler (hereafter “KL”) divergence from $\mathbb{P}$ and $\mathbb{P}^{(0)}$ equals [^19] See Rasmussen and Williams, (2006, Section A.5) for a derivation of (7).
@@
The KL divergence (7) measures the information gain from using $\mathbb{P}$ as a prior rather than $\mathbb{P}^{(0)}$. This information is purely conceptual: it does not depend on the sample $\mathcal{S}$}. It comes from knowing how to represent states as low-dimensional combinations of high-dimensional concepts. We study the instrumental value of this dimension reduction in Section 5.
@@
The KL divergence (7) equals zero when the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ of $\Sigma$ are equal and is larger when they are more spread out (see Proposition A4). [^20] If $\mathbb{P}^{(0)}$ has mean $\mu^{(0)}\in\mathbb{R}^{K}$, then (7) becomes $\mathcal{D}_{\text{KL}}(\mathbb{P}\parallel\mathbb{P}^{(0)})=\frac{1}{2}\mathopen{}\mathclose{{\left(\frac{1}{\overline{\lambda}}\lVert\mu-\mu^{(0)}\rVert-\sum_{k=1}^{K}\ln\mathopen{}\mathclose{{\left(\frac{\lambda_{k}}{\overline{\lambda}}}}\right)}}\right).$ So even if $\mu^{(0)}\not=\mu$, the KL divergence from $\mathbb{P}$ to $\mathbb{P}^{(0)}$ is non-negative and does not fall when $\lambda_{1},\ldots,\lambda_{K}$ undergo a MPS (see Proposition A4). But it is strictly larger than zero when $\mu^{(0)}\not=\mu$, even if $\lambda_{1}=\cdots=\lambda_{K}$. This is because spreading $\lambda_{1},\ldots,\lambda_{K}$ narrows the distribution $\mathbb{P}$ to a lower-dimensional subspace of $\mathbb{R}^{K}$.
For example, consider the prior variance matrix (1) derived in Section 2. This matrix has eigenvalues $\lambda_{1}=\sigma^{2}(1+\rho)$ and $\lambda_{2}=\sigma^{2}(1-\rho)$, which have mean $\overline{\lambda}=\sigma^{2}$ and become more spread out as $\rho\in[0,1)$ grows. The KL divergence
$\mathcal{D}_{\text{KL}}(\mathbb{P}\parallel\mathbb{P}^{(0)})=-\frac{1}{2}\ln(1-\rho^{2})$
from $\mathbb{P}$ to $\mathbb{P}^{(0)}$ equals zero when $\rho=0$ and grows as $\rho$ grows. If $\rho=0$, then the true prior $\mathbb{P}$ has equal variance in all directions of $\mathbb{R}^{2}$, and so the agent gains nothing from knowing the eigenvectors $v_{1}$ and $v_{2}$ of (1). The larger is $\rho$, the more concentrated is $\mathbb{P}$ around the subspace spanned by $v_{1}$, and so the more the agent gains from knowing $v_{1}$ and $v_{2}$.
Relationship to PCA
The example above illustrates the connection between our ideas and principal component analysis (hereafter “PCA”). PCA is a dimension reduction technique that projects a distribution onto its highest variance dimensions. Traditional PCA estimates these dimensions from data. In contrast, our agent derives them from his conceptual knowledge: he knows which dimensions have the highest variance before observing any data. This “pre-data PCA” allows him to collect more instrumentally valuable data; we call this benefit the “value of conceptual knowledge” and quantify it in Sections 57.
3.3 Modeling assumptions
We assume states and outcomes are jointly normally distributed under the agent’s prior, and his actions are real-valued and induce quadratic losses. This setup is common in the literature on statistical decisions (Hastie et al.,, 2009). It is also implicit in empirical economics papers that estimate linear models via Ordinary Least Squares. Our agent has a linear model (5) of the unknown state vector. If his prior is diffuse, then his optimal actions equal the estimates obtained via OLS.
We also assume the agent knows how the states covary a priori. This separates the conceptual knowledge embedded in his prior from the statistical knowledge he infers from his sample. The assumption allows us to measure the agent’s conceptual and statistical knowledge on independent scales, and to study their relative contributions to his welfare (see Sections 6 and 7).
Finally, we assume there is a correct model of the agent’s environment (i.e., a true prior variance matrix) that he can know at different “depths” (see Sections 6 and 7). This separates our paper from the literatures on model uncertainty (Chatfield,, 1995; Marinacci,, 2015) and mis-specification (Esponda and Pouzo,, 2016; Spiegler,, 2016), which study agents who do not know the correct model or use an incorrect model. Our analysis complements those literatures: rather than asking “what if the agent does not know the correct model?” we ask “what does he gain from knowing the correct model?” [^21] However, our analysis connects to robust approaches (Gilboa and Schmeidler,, 1989; Hansen and Sargent,, 2001; Klibanoff et al.,, 2005). The naïve prior is robust in that it commits minimally to any particular covariance structure; it spreads variance evenly across all dimensions. Theorem 1 can be interpreted as saying the cost of this robust approach is larger when the correct model has more structure (i.e., the eigenvalues $\lambda_{1},\ldots,\lambda_{K}$ are more spread out). In this way, we quantify the value of imposing correct structural restrictions versus maintaining robustness to mis-specification.